8-A block weighs one pound on a scale at the supermarket. When put in water it sinks. Arn underwater scale shows a weight of 1/3 of a pound. What is the buoyant force? 9-A block weighs one pound on a scale at the supermarket. When put in water it sinks. An underwater scale shows a 1/3 of a pound. What is the density of the block? 10-A block of the same material as in problem 9 weighs two pounds and is submerged in water. What weight would an underwater scale show?
Answer #1

here,

8)

the weight of block in air , Wa = 1 pound

the weight of block in water , Ww = 1/3 pound

the Buoyant force , F = Wa - Ww

F = 1 - 1/3 pounds

F = 2/3 pounds

F = 0.67 pounds

the buoyant force is 0.67 pounds